Zoe has a solution that is 75% acid and a solution that is 25% acid. How much of each should she use to obtain 400ml of a solution that is 45% acid?
Answer by robertb(5830)

0.75x + 0.25(400-x) = 0.45*400,
0.75x + 100-0.25x = 180,
0.50x = 80,
x = 160 ml of the 75% acid solution, and
400 - x = 240 ml of the 25% acid solution.